OpenStudy (anonymous):

What is the integral of { [(H-h)^2] / sqrt h } dh ?

OpenStudy (anonymous):



OpenStudy (anonymous):

ok, u have\[\frac{(H-h)^2}{\sqrt{h}}=\frac{H^2-2hH+h^2}{\sqrt{h}}=H^2h^{-1/2}-2Hh^{1/2}+h^{3/2}\]do u know what is the anti-derivative of \(h^n\) because all the terms are in this form

OpenStudy (anonymous):

yes \[(h ^{n+1}) / n+1\]

OpenStudy (anonymous):

very good :)

OpenStudy (anonymous):

thanks :)